### v1

# EMD: Explicit Motion Modeling for High-Quality Street Gaussian Splatting

Abstract
Photorealistic reconstruction of street scenes is essential for developing real-world simulators in autonomous driving. While recent methods based on 3D/4D Gaussian Splatting (GS) have demonstrated promising results, they still encounter challenges in complex street scenes due to the unpredictable motion of dynamic objects. Current methods typically decompose street scenes into static and dynamic objects, learning the Gaussians in either a supervised manner (e.g., w/ 3D bounding-box) or a self-supervised manner (e.g., w/o 3D bounding-box). However, these approaches do not effectively model the motions of dynamic objects (e.g., the motion speed of pedestrians is clearly different from that of vehicles), resulting in suboptimal scene decomposition. To address this, we propose Explicit Motion Decomposition (EMD), which models the motions of dynamic objects by introducing learnable motion embeddings to the Gaussians, enhancing the decomposition in street scenes. The proposed EMD is a plug-and-play approach applicable to various baseline methods. We also propose tailored training strategies to apply EMD to both supervised and self-supervised baselines. Through comprehensive experimentation, we illustrate the effectiveness of our approach with various established baselines. The code will be released at: https://qingpowuwu.github.io/emdgaussian.github.io/.
11footnotetext: Equal contribution.22footnotetext: Work done during internship at NIO.33footnotetext: Corresponding Author.
{strip}[Uncaptioned image]
Figure 1: Previous street Gaussian splatting methods find it challenging to accurately model the motion of dynamic objects, which leads to blurry reconstructions. With the introduction of the proposed Explicit Motion Decomposition (EMD), which improves the decomposition of dynamic object motion, the current method achieves significantly better reconstruction quality.
1 Introduction
Novel view synthesis for dynamic scenes is essential in autonomous driving, enabling a variety of applications such as simulation, testing, and validation of perception systems. Traditional simulators like CARLA (Dosovitskiy et al. 2017) and AirSim (Shah et al. 2018), while providing controlled environments, suffer from limited realism and require substantial manual effort in creating virtual scenes. Recent advances in neural rendering, particularly Neural Radiance Fields (NeRF) (Mildenhall et al. 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), have emerged as promising alternatives for photorealistic scene reconstruction. These pioneering approaches excel in capturing complex geometries and appearances through implicit or explicit neural representations, and have been extended to dynamic scenes by incorporating an additional time dimension into their representations and learning deformation networks to model non-rigid motions (Park et al. 2021a; Pumarola et al. 2020; Li et al. 2021; Tretschk et al. 2020; Park et al. 2021b; Xian et al. 2020; Xu, Alldieck, and Sminchisescu 2021; Wang et al. 2022; Cleac’h et al. 2022; Cao and Johnson 2023; Rabich, Stotko, and Klein 2023). Despite advancements, reconstructing autonomous driving scenes remains difficult due to complex multi-object dynamics, expansive environments, and varied motion patterns.
To tackle this challenge, existing works usually build upon dynamic NeRF and 3DGS frameworks, further separating autonomous driving scenes into static and dynamic components through two primary paradigms: Supervised methods utilize pre-trained models to acquire auxiliary conditions, such as segmentation masks from SAM (Kirillov et al. 2023), depth maps from DepthAnything (Yang et al. 2024a), and both depth and optical flow from Dynamo (Sun and Hariharan 2023), or 3D bounding boxes from various datasets. Representative works like StreetGaussian (Yan et al. 2024) demonstrate the effectiveness of using supervision signals for trajectory optimization and appearance modeling. In contrast, self-supervised methods achieve static-dynamic separation without explicit supervision, as shown by S3Gaussian (Huang et al. 2024), which leverages inherent motion cues. While both paradigms have shown promising results, their binary classification of scene elements as either static or dynamic overlooks the continuous spectrum of motion inherent in real-world street scenes.
To better address the different motion patterns in street scenes, we propose an Explicit Motion Decomposition (EMD) module that can be easily integrated into existing supervised and self-supervised frameworks (Fig. 2). EMD improves scene decomposition by incorporating motion-aware feature encoding and dual-scale deformation modeling. Specifically, we enhance each Gaussian primitive with learnable motion embeddings to capture its motion characteristics and design a hierarchical deformation framework that separately manages fast, global motions and slow, local deformations. This design allows for more efficient analysis of complex street scenes with varying motion speeds.
To showcase the versatility of our approach, we conduct extensive experiments on the Waymo-NOTR dataset by integrating EMD with representative methods from both paradigms: StreetGaussian for supervised and S3Gaussian for self-supervised settings. Our main contributions include:
  • We propose EMD, a new plug-and-play module that effectively addresses varying motion speeds in street scenes through explicit motion modeling.
  • We validate our method in both supervised and self-supervised settings, demonstrating consistent performance improvements across various evaluation protocols.
  • We carried out a thorough evaluation of the Waymo-NOTR dataset, showcasing substantial improvements across various settings, achieving better reconstruction quality in both full scenes (+1.81 PSNR) and vehicle-specific regions (+2.81 PSNR) compared to respective baselines.
2 Related Work
Dynamic Scene Representation.
Neural Radiance Fields (NeRF) (Mildenhall et al. 2020) revolutionized novel view synthesis by introducing volumetric scene representation using MLPs(Taud and Mas 2018), and has been significantly improved through various extensions (Barron et al. 2021, 2022; Müller et al. 2022; Liu et al. 2020). However, these methods face challenges in dynamic scene reconstruction due to the lack of temporal modeling capabilities. A common solution involves introducing time-dependent modeling through additional time conditions and deformation networks, as adopted by (Park et al. 2021a; Pumarola et al. 2020; Park et al. 2021b). Additionally, several works also improve NeRF’s reconstruction quality in both static and dynamic scenes by introducing additional supervision signals, such as semantic segmentation (Kundu et al. 2022; Zhi et al. 2021; Jain, Tancik, and Abbeel 2021) , depth maps (Xu et al. 2022; Guo et al. 2022; Deng et al. 2022a; Niemeyer et al. 2022; Deng et al. 2022b), and optical flow (Chen and Tsukada 2022; Li et al. 2023). Despite these advances, NeRF-based methods struggle with long training times and limited ability to handle complex motions. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) has demonstrated remarkable potential by representing scenes with explicit 3D Gaussian primitives, achieving both high training efficiency and superior rendering quality. Inspired by NeRF-based approaches in handling dynamic scenes, several works have extended 3DGS to model temporal changes. Specifically, (Luiten et al. 2024) proposes dynamic modeling by tracking Gaussian properties (position, orientation) while maintaining persistent appearance, while (Yang et al. 2024b) extends the 3D representation to 4D by introducing temporal Gaussian distributions. Building upon these ideas, (Yang et al. 2023b) designs dedicated deformation networks for motion modeling, (Wu et al. 2024) advances the efficiency by maintaining only canonical 3D Gaussians with HexPlane and (Bae et al. 2024) introduces embeddings to model fast and slow motions. These methods lay the groundwork for dynamic scene reconstruction in autonomous driving scenarios.
Autonomous Driving Simulation
Traditional autonomous driving simulators like AirSim (Shah et al. 2018) and CARLA (Dosovitskiy et al. 2017) require extensive manual effort for environment creation while struggling to achieve photorealistic rendering. To address this, Neural based approaches have emerged as promising solutions for simulating street scenes. Early NeRF-based methods (Ost et al. 2021; Turki, Ramanan, and Satyanarayanan 2021; Rematas et al. 2022; Huang et al. 2023) introduced neural scene representation for large-scale urban environments, followed by improvements in efficiency and scalability (Tancik et al. 2022; Turki et al. 2023; Lu et al. 2023; Liu et al. 2023). For dynamic urban scene reconstruction, NSG (Ost et al. 2021) introduced supervised scene decomposition using learned scene graphs and latent object representations. This supervised approach has since become a new paradigm, adopted by MARS (Wu et al. 2023), NeuRAD (Tonderski et al. 2023) and Multi-Level Neural Scene Graphs (Fischer et al. 2024). With the advent of 3D Gaussian Splatting, DrivingGaussian (Zhou et al. 2024b) also adopted this supervised paradigm and developed hierarchical scene representations combining dynamic object graphs with incrementally updated static elements, while StreetGaussian (Yan et al. 2024) enhanced this framework with trajectory optimization and appearance modeling using 4D spherical harmonics. Recent work HUGS (Zhou et al. 2024a) further advances this line of research by incorporating physical constraints into the joint optimization of geometry, appearance, and semantics. To eliminate the need for expensive supervision, VDG (Li et al. 2024) introduced pose-free reconstruction by integrating self-supervised visual odometry for pose estimation and depth initialization, while employing motion mask supervision for static-dynamic scene decomposition. Inspired by D2NeRF’s (Wu et al. 2022) self-supervised decomposition in general dynamic scenes, SUDS (Turki et al. 2023) introduced this paradigm to autonomous driving by leveraging optical flow guidance, followed by EmerNeRF (Yang et al. 2023a) proposing a flow-free approach with hash-grid based scene organization. Despite the dominance of NeRF-based methods, this self-supervised paradigm was later extended to 3D Gaussian Splatting by several works. PVG (Chen et al. 2023) pioneered this extension by introducing periodic vibration-based temporal dynamics for unified representation of both static and dynamic elements, followed by S3Gaussian (Huang et al. 2024).
However, current methods typically employ binary static-dynamic classification, which oversimplifies the complex motion patterns in street scenes. This limitation becomes particularly evident when reconstructing objects with varying velocities, as shown in our evaluations on the Waymo-NOTR dataset. To address this challenge, we propose Explicit Motion Decomposition (EMD), a plug-and-play approach that better captures the continuous spectrum of motion in street scenes. Before presenting our method, we first review the fundamentals of 3D and 4D Gaussian Splatting.
Figure 2: Overview of our Explicit Motion Decomposition (EMD) framework. Given input Gaussian primitives, our method processes them through two main components: (a) Motion-aware Feature Encoding, which combines spatial, temporal, and Gaussian-specific information to capture motion characteristics; and (b) Dual-scale Deformation Framework, which hierarchically models fast global motions and slow local deformations. The framework can be seamlessly integrated into both supervised (StreetGaussian) and self-supervised (S3Gaussian) approaches through our proposed integration strategies.
3 Preliminaries
3.1 3D Gaussian Splatting
3D Gaussian Splatting (3D-GS) (Kerbl et al. 2023) proposes an explicit rendering approach to represent 3D scenes through a collection of 3D Gaussian primitives $\mathbb{G}=\left\{\left(\mu_{k},\Sigma_{k},\alpha_{k},\mathbf{c}_{k}\right)\right\}_{k=1}^{K}$, where $K$ is the total number of Gaussians. Each Gaussian primitive represents a probabilistic distribution of density in 3D space, defined by its probability density function:
$G_{k}(x)=e^{-\frac{1}{2}(x-\mu_{k})^{T}\Sigma_{k}^{-1}(x-\mu_{k})},$ (1)
where $x\in\mathbb{R}^{3}$ represents any point in the 3D world space, $\mu_{k}\in\mathbb{R}^{3}$ and $\Sigma_{k}\in\mathbb{R}^{3\times 3}$ denote the mean position and covariance matrix in world space, respectively, where the covariance matrix determines the shape and orientation of the Gaussian, $\alpha_{k}\in[0,1]$ is the opacity, and $\mathbf{c}_{k}$ encodes the view-dependent color information.
Rendering Process
For rendering, each 3D Gaussian is projected onto the image plane, where the 3D mean $\mu_{k}$ is transformed to 2D mean $\mu_{k}^{2D}$, and the world space covariance matrix $\Sigma_{k}$ is transformed to screen space as $\Sigma_{k}^{\prime}=JW\Sigma_{k}W^{T}J^{T}$, with $W$ and $J$ being the viewing transformation and projective transformation Jacobian matrices, respectively. The final pixel color at screen space position $x$ is computed through front-to-back alpha compositing:
$C(x)=\sum_{k\in\mathcal{N}(\mathbf{x})}\mathbf{c}_{k}\alpha_{k}(x)\prod_{j=1}^{k-1}\left(1-\alpha_{j}(x)\right),$ (2)
where $\alpha_{k}(x)=\alpha_{k}\exp\left(-\frac{1}{2}\left(x-\mu_{k}^{2D}\right)^{T}{\Sigma_{k}^{\prime}}^{-1}\left(x-\mu_{k}^{2D}\right)\right)$ represents the opacity contribution of the $k$-th Gaussian at pixel $x$, computed as the product of the Gaussian’s base opacity $\alpha_{k}$ and its projected 2D Gaussian evaluation. $\mathcal{N}(\mathbf{x})$ represents the set of indices of Gaussians intersecting pixel $x\in\mathbb{R}^{2}$.
In practice, for each Gaussian $k$, we parameterize its covariance matrix $\Sigma_{k}$ using rotation quaternion $\mathbf{q}_{k}$ and scaling vector $\mathbf{s}_{k}$ as:
$\Sigma_{k}=R(\mathbf{q}_{k})S(\mathbf{s}_{k})S(\mathbf{s}_{k})^{T}R(\mathbf{q}_{k})^{T},$ (3)
where $R(\mathbf{q}_{k})$ is the rotation matrix defined by quaternion $\mathbf{q}_{k}$, and $S(\mathbf{s}_{k})$ is the diagonal scaling matrix defined by scaling vector $\mathbf{s}_{k}$. The view-dependent color $\mathbf{c}_{k}$ is encoded using spherical harmonics (SH) coefficients:
$\mathbf{c}_{k}(\mathbf{d})=\sum_{l=0}^{L}\sum_{m=-l}^{l}k_{l,m}^{(k)}Y_{l,m}(\mathbf{d}),$ (4)
where $\mathbf{d}$ is the viewing direction, $Y_{l,m}$ are the spherical harmonic basis functions, and $k_{l,m}^{(k)}$ are the corresponding coefficients for the $k$-th Gaussian.
3.2 4D Gaussian Splatting
4D Gaussian Splatting extends the static 3D-GS framework to handle dynamic scenes by incorporating temporal information. For a dynamic scene captured at different timestamps $t\in[0,T]$, each Gaussian primitive is now characterized by time-varying parameters: $\mathbb{G}(t)=\left\{\left(\mu_{k}(t),\Sigma_{k}(t),\alpha_{k}(t),\mathbf{c}_{k}(t)\right)\right\}_{k=1}^{K}$. The probability density function of each Gaussian at time $t$ becomes:
$G_{k}(x,t)=e^{-\frac{1}{2}(x-\mu_{k}(t))^{T}\Sigma_{k}(t)^{-1}(x-\mu_{k}(t))},$ (5)
The temporal evolution of these parameters is typically modeled through one of two main approaches:
Deformation-based Approach
This approach models temporal changes by applying a deformation field to the base Gaussians. For each timestamp $t$, the position of each Gaussian is updated as:
$\mu_{k}(t)=\mu_{k}(0)+\Delta\mu_{k}(t),$ (6)
where $\mu_{k}(0)$ is the initial position and $\Delta\mu_{k}(t)$ is the displacement predicted by a deformation network. Similarly, other parameters such as rotation, scaling, and color can be modeled as temporal offsets from their initial states.
Trajectory-based Approach
Alternatively, the temporal evolution can be represented by explicitly modeling the continuous trajectory of each parameter. For instance, the position trajectory can be parameterized using a set of N control points $\{\mathbf{p}_{i}\}_{i=1}^{N}$ and basis functions $\{\phi_{i}(t)\}_{i=1}^{N}$:
$\mu_{k}(t)=\sum_{i=1}^{N}\mathbf{p}_{i}\phi_{i}(t),$ (7)
where the basis functions $\phi_{i}(t)$ can be B-splines or other temporal interpolation functions. The rendering process remains similar to static 3D-GS, but now Gaussian parameters are evaluated at the specific timestamp $t$ before projection and compositing.
4 Methodology
4D Gaussian Splatting has shown promising results in dynamic scene reconstruction. However, modeling dynamic street scenes remains challenging due to diverse motion patterns. To better handle the varying motion patterns in street scenes, we propose Explicit Motion Decomposition (EMD), a plug-and-play module that can be seamlessly integrated into existing 4D Gaussian-based frameworks to enhance their capability in handling dynamic scenarios, as illustrated in Fig. 2.
4.1 Problem Formulation
Given a set of static 3D Gaussian primitives $\mathbb{G}=\{(\mu_{k},\mathbf{s}_{k},\mathbf{q}_{k},\alpha_{k},\mathbf{c}_{k})\}_{k=1}^{K}$ and a timestamp $t$, our goal is to learn a deformation field $\mathcal{D}$ that maps each Gaussian’s parameters from their canonical states to their corresponding deformed states at time $t$. For notational simplicity, we omit the Gaussian index $k$ in the following formulation:
$\{\mu_{t},\mathbf{s}_{t},\mathbf{q}_{t},\alpha_{t},\mathbf{c}_{t}\}=\mathcal{D}(\{\mu,\mathbf{s},\mathbf{q},\alpha,\mathbf{c}\},t),$ (8)
To effectively handle the diverse motion patterns in street scenes, especially the distinct movements between vehicles and pedestrians, we propose a motion-aware deformation module that processes input Gaussian parameters through two key components: motion-aware feature encoding and dual-scale deformation prediction.
Motion-aware Feature Encoding
Different types of objects in street scenes exhibit distinct motion characteristics. To capture these varied patterns, we first encode the input Gaussian parameters into a comprehensive feature space that combines spatial, temporal, and Gaussian-specific information:
$\mathbf{F}_{aggr}(\mu,t)=[\mathbf{F}_{pos}(\mu),\mathbf{F}_{temp}(t),\mathbf{F}_{gauss}],$ (9)
The spatial component employs multi-frequency positional encoding:
$\mathbf{F}_{pos}(\mu)=[\mu,\{\sin(2^{i}\pi\mu),\cos(2^{i}\pi\mu)\}_{i=0}^{P-1}],$ (10)
where $P$ is the number of frequency bands, enabling the network to capture both fine geometric details and global structures. For temporal information, we design an adaptive temporal embedding function:
$\mathbf{F}_{temp}(t)=\mathcal{T}(t,N(i))=\text{Interp}(\mathbf{W},t,N(i)),$ (11)
where $\mathbf{W}\in\mathbb{R}^{N_{max}\times D}$ is a learnable embedding matrix that captures motion patterns, and $N(i)$ progressively increases from $N_{min}$ to $N_{max}$ temporal samples during training iteration $i$, allowing the model to gradually capture finer temporal dynamics. For Gaussian-specific features, we assign a learnable latent embedding $\mathbf{z}_{k}\in\mathbb{R}^{D}$ to each Gaussian $k$, where $\mathbf{F}_{gauss}=\mathbf{z}_{k}$, enabling the model to learn and represent individual motion characteristics.
Table 1: Comparative performance of our framework and baseline approaches on the Waymo-NOTR dataset. The best performances are highlighted in bold, and the second-best are indicated with underlining. ↑ indicates higher is better, while ↓ indicates lower is better.
Scene Reconstruction Novel View Synthesis
Dataset Methods Full Image Vehicle Full Image Vehicle
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑
D32 EmerNeRF 28.16 0.806 0.228 24.32 0.682 25.14 0.747 0.313 23.49 0.660
3DGS 28.47 0.876 0.136 23.26 0.716 25.14 0.813 0.165 20.48 0.753
MARS 28.24 0.866 0.252 23.37 0.701 26.61 0.796 0.305 22.21 0.697
S3Gaussian 30.69 0.900 0.121 26.23 0.804 26.62 0.824 0.159 22.61 0.681
S3Gaussian+Ours 32.50 0.933 0.082 29.04 0.879 26.55 0.833 0.126 23.39 0.703
Dual-scale Deformation Framework
Given the diverse motion patterns in street scenes, ranging from large vehicular movements to subtle pedestrian motions, we design a hierarchical deformation framework that can effectively handle both scales of motion:
$\begin{split}\mathcal{D}(\mu,t)=&\mathcal{D}_{coarse}(\mathbf{F}_{aggr}(\mu,t))\\ &+\mathcal{D}_{fine}(\mathbf{F}_{aggr}(\mu+\Delta\mu_{coarse},t)),\end{split}$ (12)
The final deformed parameters combine both coarse and fine scale predictions:
$\mu_{t}$ $=\mu+\Delta\mu_{coarse}+\Delta\mu_{fine}$ (13)
$\mathbf{s}_{t}$ $=\mathbf{s}+\Delta\mathbf{s}_{coarse}+\Delta\mathbf{s}_{fine}$
$\mathbf{q}_{t}$ $=\mathbf{q}\otimes\Delta\mathbf{q}_{coarse}\otimes\Delta\mathbf{q}_{fine},$
where $\mathcal{D}_{coarse}$ focuses on modeling large-scale motions such as vehicle translations, while $\mathcal{D}_{fine}$ captures local deformations like articulated movements. Similar to position updates, other Gaussian parameters including opacity $\alpha_{t}$ and spherical harmonics coefficients $\mathbf{c}_{t}$ are also updated through this dual-scale framework.
Figure 3: Qualitative comparison on the self-supervised setting between S3Gaussian and S3Gaussian+ours EMD. We also visualize the error maps between the rendered images and ground truth to provide further insights.
4.2 Integration with Existing Frameworks
As mentioned in the introduction, current approaches for street scene reconstruction generally fall into supervised and self-supervised paradigms. Having formalized our EMD framework, we now demonstrate its integration into representative methods from both paradigms: StreetGaussian for supervised learning and S3Gaussian for self-supervised learning.
Self-supervised Integration: S3Gaussian
For self-supervised scenarios, we enhance S3Gaussian’s architecture with our motion-aware features. While S3Gaussian originally employs a Multi-head Gaussian Decoder for deformation prediction, we augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and restructure its decoder into our dual-scale framework. Specifically, both coarse and fine stages predict deformations in position ($\Delta\mu$), opacity ($\Delta\alpha$), and spherical harmonics coefficients ($\Delta\mathbf{c}$), enabling more precise motion modeling through hierarchical refinement.
Supervised Integration: StreetGaussian
For supervised settings, StreetGaussian provides a framework that represents dynamic objects through tracked vehicle poses and object-specific Gaussians. Each object is characterized by tracked poses $\{R_{t},T_{t}\}_{t=1}^{N_{t}}$ that transform object Gaussians from local coordinates ($\mu_{o},R_{o}$) to world coordinates ($\mu_{w},R_{w}$). To enhance its motion modeling capability, we also augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and apply temporal embedding only within each object’s tracked duration, effectively capturing object-specific temporal dynamics.
To address the challenge of noisy tracked poses while maintaining their geometric meaning, we incorporate our dual-scale framework into their pose optimization:
$R^{\prime}_{t}$ $=R_{t}(\Delta R_{t}^{c}+\Delta R_{t}^{f})$ (14)
$T^{\prime}_{t}$ $=T_{t}+(\Delta T_{t}^{c}+\Delta T_{t}^{f}),$
where $\Delta R_{t}^{c},\Delta T_{t}^{c}$ handle large pose corrections and $\Delta R_{t}^{f},\Delta T_{t}^{f}$ capture subtle adjustments. This decomposition allows the model to effectively correct tracking errors while preserving the physical meaning of the tracked poses. Similarly, we apply the dual-scale framework to appearance modeling, where spherical harmonics coefficients are refined through both coarse and fine stages, enabling more accurate dynamic appearance representation.
We adopt the same loss function as S3Gaussian and StreetGaussian to train the entire pipeline. For further training details, please refer to the supplementary materials.
5 Experiments
Figure 4: Motion deformation comparison between S3Gaussian + Ours and S3Gaussian. Please zoom in for more details.
In this section, we first describe the datasets and evaluation metrics used in our experiments (Sec. 5.1). Then, we demonstrate the experimental results of our approach under both self-supervised and supervised settings (Sec. 5.2). Finally, we present comprehensive ablation studies to validate the effectiveness of individual components in our framework (Sec. 5.3).
5.1 Datasets and Metrics
Dataset.
We conduct extensive experiments on the Waymo Open Dataset (Sun et al. 2020) to comprehensively evaluate our method. To validate the effectiveness of our plug-and-play module, we select representative state-of-the-art methods from both supervised and self-supervised paradigms: StreetGaussian (Yan et al. 2024) and S3Gaussian (Huang et al. 2024). For self-supervised evaluation, we use the dynamic32 (D32) split introduced by EmerNeRF (Yang et al. 2023a), containing 32 sequences with vehicle motion, where each sequence consists of approximately 50-100 frames captured by three cameras under various conditions. For supervised evaluation, we follow the data split protocol established in StreetGaussian (Yan et al. 2024) to enable direct comparison.
Evaluation Metrics.
We employ comprehensive metrics to evaluate both reconstruction quality and novel view synthesis capability. For scene reconstruction, we use PSNR and SSIM to evaluate both full-scene and vehicle-specific reconstruction quality. Additionally, we compute LPIPS for perceptual quality assessment. Following previous protocols, we evaluate novel view synthesis on every 10th frame for self-supervised settings and every 4th frame for supervised settings.
Figure 5: Qualitative ablation study results across three camera views from the Waymo dataset. (a) Our complete model achieves sharp and consistent reconstruction. (b) Removing Gaussian embedding $\mathcal{F}_{gauss}$ leads to blurry object boundaries. (c) Without temporal embedding $\mathcal{F}_{temp}$, results show motion artifacts. (d) Without coarse deformation $\mathcal{D}_{coarse}$, geometric consistency is lost. (e) Absence of fine deformation $\mathcal{D}_{fine}$ causes detail degradation.
5.2 Main Results
Self-supervised Performance
Tab. 1 presents the comparative results on the D32 split, where no 3D bounding box annotations were used. Our method significantly outperforms previous self-supervised approaches, achieving notable improvements in both full-scene and object-specific metrics. Specifically, we observe a substantial increase in PSNR for the full scene (32.50 vs. 30.69) and for vehicle-specific metrics (PSNR: 29.04 vs. 26.23) when compared to S3Gaussian. These results demonstrate the enhanced ability of our method to accurately model complex street scenes without explicit 3D box annotations. For novel view synthesis, our method continues to perform competitively, achieving a PSNR of 26.55, which reflects its robust generalization to previously unseen viewpoints.
In addition, we present a side-by-side visualization comparison between S3Gaussian and S3Gaussian+EMD in Fig. 3. The error maps, which compare the ground truth to the rendered results, clearly demonstrate that S3Gaussian+EMD outperforms S3Gaussian in modeling dynamic objects with varying motion speeds. S3Gaussian implicitly models dynamic vehicles, but it fails to capture changes in speed during motion and the differences in movement between vehicles, leading to blurred reconstructions. on the contrary, our method captures the distinct motion characteristics of different dynamic objects, leading to more accurate and consistent scene reconstructions. This emphasizes the effectiveness of Explicit Motion Decomposition (EMD) in modeling the motion of dynamic objects, improving the overall decomposition and photorealistic rendering of street scenes.
We also present a visual comparison of motion between S3Gaussian and S3Gaussian+EMD. As shown in Fig. 4, the proposed dual-scale deformation network generates a coarse deformation to model slower motion and larger-scale geometry and a fine deformation to capture faster motion and finer geometric details in the scene. With the incorporation of EMD, S3Gaussian can capture detailed features of dynamic vehicles, including the car brand logo, as demonstrated in the figure. In contrast, S3Gaussian treats the entire moving car as a single dynamic object, failing to produce clear synthesis results for the dynamic vehicles.
Supervised Performance
Table 2: Comparative performance of our framework and baseline approaches on the StreetGaussian dataset. The best performances are highlighted in bold. ↑ indicates higher is better, while ↓ indicates lower is better.
Methods Full Image Vehicle
PSNR↑ SSIM↑ LPIPS↓ PSNR↑
3DGS 29.64 0.918 0.117 21.25
NSG 28.31 0.862 0.346 24.32
MARS 29.75 0.886 0.264 26.54
EmerNeRF 30.87 0.905 0.133 21.67
StreetGaussian 34.61 0.938 0.079 30.23
StreetGaussian + Ours 36.03 0.949 0.067 29.81
To demonstrate the versatility of our framework, we also evaluate it in the supervised setting using the same scenes as StreetGaussian (Yan et al. 2024) (Tab. 2). When incorporating 3D bounding box supervision through our adaptive training scheme, our method achieves superior performance in full scene reconstruction, improving PSNR by 1.42dB (36.03 vs. 34.61), SSIM by 1.1% (0.949 vs. 0.938), and LPIPS by 15.2% (0.067 vs. 0.079). These results demonstrate that our approach effectively enhances the overall scene reconstruction quality while maintaining competitive performance in dynamic object modeling. Please refer to the supplementary for visualization on the supervised setting.
5.3 Ablation Studies
Table 3: Ablation study on the D32 dataset showing the impact of different components in our framework. All variants are evaluated using the self-supervised setting. The best performances are highlighted in bold. ↑ indicates higher is better, while ↓ indicates lower is better.
Variant Full Image Vehicle
PSNR↑ SSIM↑ LPIPS↓ PSNR↑
Full Model 32.50 0.933 0.082 29.04
w/o Gaussian Embedding 32.21 0.928 0.089 28.80
w/o Temporal Embedding 32.23 0.922 0.091 28.08
w/o Coarse Deformation 29.40 0.890 0.146 24.54
w/o Fine Deformation 32.45 0.931 0.118 28.80
To assess the contribution of each component in our framework, we perform comprehensive ablation studies, as shown in Tab. 3 and Fig. 5. The results reveal the critical role of the Gaussian embedding, as its removal leads to the performance drop, with a reduction of 0.29 PSNR (32.21 vs. 32.50). This indicates that the Gaussian embedding is essential for effectively capturing the motion characteristics for each dynamic gaussian. The temporal embedding also plays a crucial role, with its absence leading to a 0.27 PSNR drop (32.23 vs. 32.50), underscoring its importance in modeling the temporal variation of object motion over time.
Both the coarse and fine deformation components are integral to the final performance, with their removal leading to considerable performance degradation. Specifically, excluding the coarse deformation component causes a significant 3.10 PSNR drop (29.40 vs. 32.50), suggesting that the coarse adjustments are vital for maintaining the overall scene structure. On the other hand, removing the fine deformation component results in a 0.036 LPIPS increase (0.118 vs. 0.082), implying that fine deformation refines the details and its absence worsens the perceptual quality. The ablation study demonstrates that our proposed motion-aware feature encoding and dual-scale deformation effectively model dynamic objects with varying motion speeds, which is crucial for enhancing the reconstruction quality of existing street Gaussians.
6 Limitation
Although EMD effectively addresses the challenge of modeling dynamic objects with varying speeds by incorporating learnable embeddings, some limitations remain. Existing street Gaussian methods do not account for environmental lighting, yet lighting effects play a crucial role in the quality of reconstructions under different lighting conditions. In future work, we plan to explore the possibility of developing a plug-and-play technique to enhance lighting effects in existing methods.
7 Conclusion
In this paper, we present EMD, a plug-and-play module that effectively handles varying motion speeds in street scene reconstruction. By introducing motion-aware feature encoding and dual-scale deformation modeling, our approach successfully captures the continuous spectrum of motion patterns inherent in real-world scenarios. Comprehensive experiments on the Waymo Open Dataset demonstrate that EMD significantly improves reconstruction quality when integrated with both supervised and self-supervised frameworks. Our method substantially improves full scene reconstruction (+1.81 PSNR) and vehicle-specific regions (+2.81 PSNR), validating the effectiveness of explicit motion modeling. We believe our work opens up new possibilities for high-quality dynamic scene reconstruction in autonomous driving applications, and the plug-and-play nature of our approach makes it readily applicable to future developments in neural scene representations.

### v2

@@
 To address this, we propose Explicit Motion Decomposition (EMD), which models the motions of dynamic objects by introducing learnable motion embeddings to the Gaussians, enhancing the decomposition in street scenes.
-The proposed EMD is a plug-and-play approach applicable to various baseline methods.
-We also propose tailored training strategies to apply EMD to both supervised and self-supervised baselines.
-Through comprehensive experimentation, we illustrate the effectiveness of our approach with various established baselines.
-The code will be released at: https://qingpowuwu.github.io/emdgaussian.github.io/.
+The proposed plug-and-play EMD module compensates for the lack of motion modeling in self-supervised street Gaussian splatting methods.
+We also introduce tailored training strategies to extend EMD to supervised approaches.
+Comprehensive experiments demonstrate the effectiveness of our method, achieving state-of-the-art novel view synthesis performance in self-supervised settings.
+The code is available at: https://qingpowuwu.github.io/emd.
+{strip}[Uncaptioned image]
+Figure 1: Previous street Gaussian splatting methods find it challenging to accurately model the motion patterns of dynamic objects, which leads to blurry reconstructions.
+With the introduction of the proposed Explicit Motion Decomposition (EMD), which compensates for the modeling of dynamic object motion, achieving the state-of-the-art reconstruction quality.
 11footnotetext: Equal contribution.22footnotetext: Work done during internship at NIO.33footnotetext: Corresponding author.
-{strip}[Uncaptioned image]
-Figure 1: Previous street Gaussian splatting methods find it challenging to accurately model the motion of dynamic objects, which leads to blurry reconstructions.
-With the introduction of the proposed Explicit Motion Decomposition (EMD), which improves the decomposition of dynamic object motion, the current method achieves significantly better reconstruction quality.
 1 Introduction
-Novel view synthesis for dynamic scenes is essential in autonomous driving, enabling a variety of applications such as simulation, testing, and validation of perception systems.
-Traditional simulators like CARLA (Dosovitskiy et al.
-2017) and AirSim (Shah et al.
-2018), while providing controlled environments, suffer from limited realism and require substantial manual effort in creating virtual scenes.
-Recent advances in neural rendering, particularly Neural Radiance Fields (NeRF) (Mildenhall et al.
-2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al.
-2023), have emerged as promising alternatives for photorealistic scene reconstruction.
-These pioneering approaches excel in capturing complex geometries and appearances through implicit or explicit neural representations, and have been extended to dynamic scenes by incorporating an additional time dimension into their representations and learning deformation networks to model non-rigid motions (Park et al.
-2021a; Pumarola et al.
-2020; Li et al.
-2021; Tretschk et al.
-2020; Park et al.
-2021b; Xian et al.
-2020; Xu, Alldieck, and Sminchisescu 2021; Wang et al.
-2022; Cleac’h et al.
-2022; Cao and Johnson 2023; Rabich, Stotko, and Klein 2023).
-Despite advancements, reconstructing autonomous driving scenes remains difficult due to complex multi-object dynamics, expansive environments, and varied motion patterns.
-To tackle this challenge, existing works usually build upon dynamic NeRF and 3DGS frameworks, further separating autonomous driving scenes into static and dynamic components through two primary paradigms: Supervised methods utilize pre-trained models to acquire auxiliary conditions, such as segmentation masks from SAM (Kirillov et al.
-2023), depth maps from DepthAnything (Yang et al.
-2024a), and both depth and optical flow from Dynamo (Sun and Hariharan 2023), or 3D bounding boxes from various datasets.
-Representative works like StreetGaussian (Yan et al.
-2024) demonstrate the effectiveness of using supervision signals for trajectory optimization and appearance modeling.
-In contrast, self-supervised methods achieve static-dynamic separation without explicit supervision, as shown by S3Gaussian (Huang et al.
-2024), which leverages inherent motion cues.
-While both paradigms have shown promising results, their binary classification of scene elements as either static or dynamic overlooks the continuous spectrum of motion inherent in real-world street scenes.
-To better address the different motion patterns in street scenes, we propose an Explicit Motion Decomposition (EMD) module that can be easily integrated into existing supervised and self-supervised frameworks (Fig.
+Novel view synthesis for dynamic street scenes is essential in closed-loop autonomous driving.
+Recent advances in neural rendering, particularly Neural Radiance Fields (NeRF) [31] and 3D Gaussian Splatting (3DGS) [19], have emerged as promising scene reconstruction methods.
+These pioneering approaches excel in capturing complex geometries and appearances through implicit or explicit neural representations and have been extended to dynamic scenes  [36, 39, 23, 48, 37, 56, 59, 51, 7, 3, 40, 53, 52].
+Despite advancements, reconstructing autonomous driving scenes remains difficult due to complex multi-object dynamics, complex environments, and varied motion patterns.
+To tackle this challenge, existing works based on dynamic NeRF and 3DGS frameworks separate autonomous driving scenes into static and dynamic components through two primary paradigms: Supervised methods acquire priors for dynamic objects, such as segmentation masks [20], optical flow [45], or 3D bounding boxes from object tracking.
+Representative works like StreetGaussian [60] and OminiRe [6] demonstrate the effectiveness of using supervision signals for appearance modeling of dynamic objects.
+In contrast, self-supervised methods achieve static-dynamic separation without explicit supervision, as shown by S3Gaussian [16] and DeSiRe-GS [38], which leverage inherent motion cues to optimize a 4D street representation.
+While both paradigms yield promising results, supervised methods adopt a binary classification of scene elements as either static or dynamic, which overlooks the continuous spectrum of motion inherent in street scenes.
+Self-supervised methods optimize the entire scene holistically but neglect the varied motion speeds among objects—for example, pedestrians generally move much slower than vehicles.
+Existing supervised methods [60, 6] mitigate dynamic errors by optimizing 3D bounding boxes, yet self-supervised approaches still lack an effective motion modeling mechanism.
+To address the different motion patterns in street scenes, we propose an Explicit Motion Decomposition (EMD) module that can be easily integrated into existing self-supervised frameworks (Fig.
 2).
@@
 This design allows for more efficient analysis of complex street scenes with varying motion speeds.
-To showcase the versatility of our approach, we conduct extensive experiments on the Waymo-NOTR dataset by integrating EMD with representative methods from both paradigms: StreetGaussian for supervised and S3Gaussian for self-supervised settings.
+We conduct extensive experiments on the Waymo and KITTI datasets by integrating EMD with representative self-supervised methods: S3Gaussian and DeSiRe-GS.
+In addition, EMD can be seamlessly extended to supervised methods.
 Our main contributions include:
-  • We propose EMD, a new plug-and-play module that effectively addresses varying motion speeds in street scenes through explicit motion modeling.
-  • We validate our method in both supervised and self-supervised settings, demonstrating consistent performance improvements across various evaluation protocols.
-  • We carried out a thorough evaluation of the Waymo-NOTR dataset, showcasing substantial improvements across various settings, achieving better reconstruction quality in both full scenes (+1.81 PSNR) and vehicle-specific regions (+2.81 PSNR) compared to respective baselines.
+  • We propose EMD, the first plug-and-play module that effectively addresses varying motion speeds in street scenes through explicit motion modeling.
+  • We introduce tailored training strategies for self-supervised street Gaussian splatting methods and further extend EMD to supervised settings.
+  • Comprehensive experiments on Waymo and KITTI datasets demonstrate previous methods with EMD exhibit better reconstruction quality, achieving state-of-the-art (SOTA) performance in self-supervised settings.
 2 Related Work
 Dynamic Scene Representation.
-Neural Radiance Fields (NeRF) (Mildenhall et al.
-2020) revolutionized novel view synthesis by introducing volumetric scene representation using MLPs(Taud and Mas 2018), and has been significantly improved through various extensions (Barron et al.
-2021, 2022; Müller et al.
-2022; Liu et al.
-2020).
-However, these methods face challenges in dynamic scene reconstruction due to the lack of temporal modeling capabilities.
-A common solution involves introducing time-dependent modeling through additional time conditions and deformation networks, as adopted by (Park et al.
-2021a; Pumarola et al.
-2020; Park et al.
-2021b).
-Additionally, several works also improve NeRF’s reconstruction quality in both static and dynamic scenes by introducing additional supervision signals, such as semantic segmentation (Kundu et al.
-2022; Zhi et al.
-2021; Jain, Tancik, and Abbeel 2021) , depth maps (Xu et al.
-2022; Guo et al.
-2022; Deng et al.
-2022a; Niemeyer et al.
-2022; Deng et al.
-2022b), and optical flow (Chen and Tsukada 2022; Li et al.
-2023).
+Neural Radiance Fields (NeRF) [31] revolutionizes novel view synthesis by introducing volumetric scene representation and has been improved through various extensions [1, 2, 32, 26].
+However, these methods are restricted to static scenes.
+A common solution involves introducing additional time conditions, as adopted by [36, 39, 37].
+By introducing additional supervision signals, several works [21, 66, 18, 58, 14, 9, 33, 8, 4, 24] improve the rendering quality in both static and dynamic scenes.
 Despite these advances, NeRF-based methods struggle with long training times and limited ability to handle complex motions.
-Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al.
-2023) has demonstrated remarkable potential by representing scenes with explicit 3D Gaussian primitives, achieving both high training efficiency and superior rendering quality.
-Inspired by NeRF-based approaches in handling dynamic scenes, several works have extended 3DGS to model temporal changes.
-Specifically, (Luiten et al.
-2024) proposes dynamic modeling by tracking Gaussian properties (position, orientation) while maintaining persistent appearance, while (Yang et al.
-2024b) extends the 3D representation to 4D by introducing temporal Gaussian distributions.
-Building upon these ideas, (Yang et al.
-2023b) designs dedicated deformation networks for motion modeling, (Wu et al.
-2024) advances the efficiency by maintaining only canonical 3D Gaussians with HexPlane and (Bae et al.
-2024) introduces embeddings to model fast and slow motions.
+Recently, 3D Gaussian Splatting (3DGS) [19] has achieved both high efficiency and superior rendering quality by representing scenes with explicit 3D Gaussian primitives.
+Several works [30, 64, 62, 54] have extended 3DGS to incorporate temporal information, enabling the 4D scene representation.
 These methods lay the groundwork for dynamic scene reconstruction in autonomous driving scenarios.
-Autonomous Driving Simulation
-Traditional autonomous driving simulators like AirSim (Shah et al.
-2018) and CARLA (Dosovitskiy et al.
-2017) require extensive manual effort for environment creation while struggling to achieve photorealistic rendering.
-To address this, Neural based approaches have emerged as promising solutions for simulating street scenes.
-Early NeRF-based methods (Ost et al.
-2021; Turki, Ramanan, and Satyanarayanan 2021; Rematas et al.
-2022; Huang et al.
-2023) introduced neural scene representation for large-scale urban environments, followed by improvements in efficiency and scalability (Tancik et al.
-2022; Turki et al.
-2023; Lu et al.
-2023; Liu et al.
-2023).
-For dynamic urban scene reconstruction, NSG (Ost et al.
-2021) introduced supervised scene decomposition using learned scene graphs and latent object representations.
-This supervised approach has since become a new paradigm, adopted by MARS (Wu et al.
-2023), NeuRAD (Tonderski et al.
-2023) and Multi-Level Neural Scene Graphs (Fischer et al.
-2024).
-With the advent of 3D Gaussian Splatting, DrivingGaussian (Zhou et al.
-2024b) also adopted this supervised paradigm and developed hierarchical scene representations combining dynamic object graphs with incrementally updated static elements, while StreetGaussian (Yan et al.
-2024) enhanced this framework with trajectory optimization and appearance modeling using 4D spherical harmonics.
-Recent work HUGS (Zhou et al.
-2024a) further advances this line of research by incorporating physical constraints into the joint optimization of geometry, appearance, and semantics.
-To eliminate the need for expensive supervision, VDG (Li et al.
-2024) introduced pose-free reconstruction by integrating self-supervised visual odometry for pose estimation and depth initialization, while employing motion mask supervision for static-dynamic scene decomposition.
-Inspired by D2NeRF’s (Wu et al.
-2022) self-supervised decomposition in general dynamic scenes, SUDS (Turki et al.
-2023) introduced this paradigm to autonomous driving by leveraging optical flow guidance, followed by EmerNeRF (Yang et al.
-2023a) proposing a flow-free approach with hash-grid based scene organization.
-Despite the dominance of NeRF-based methods, this self-supervised paradigm was later extended to 3D Gaussian Splatting by several works.
-PVG (Chen et al.
-2023) pioneered this extension by introducing periodic vibration-based temporal dynamics for unified representation of both static and dynamic elements, followed by S3Gaussian (Huang et al.
-2024).
-However, current methods typically employ binary static-dynamic classification, which oversimplifies the complex motion patterns in street scenes.
-This limitation becomes particularly evident when reconstructing objects with varying velocities, as shown in our evaluations on the Waymo-NOTR dataset.
-To address this challenge, we propose Explicit Motion Decomposition (EMD), a plug-and-play approach that better captures the continuous spectrum of motion in street scenes.
-Before presenting our method, we first review the fundamentals of 3D and 4D Gaussian Splatting.
+Autonomous Driving Simulation.
+Traditional autonomous driving simulators like AirSim [43] and CARLA [10] require extensive manual effort for environment creation while struggling to achieve photorealistic rendering.
+To address this approaches based on neural fields have emerged as promising solutions for simulating street scenes.
+Early NeRF-based methods [34, 49, 41, 17] introduce neural scene representation for large-scale urban environments, followed by improvements in efficiency and scalability [46, 50, 28, 25].
+For dynamic urban scene reconstruction, several methods [34, 55, 47, 11] introduce supervised scene decomposition using learned scene graphs and latent object representations.
+With the advent of 3DGS, DrivingGaussian [68], StreetGaussian [60], HUGS [67] and OmniRe [6] also adopt this supervised paradigm and develop hierarchical scene representations combining dynamic object graphs with incrementally updated static elements.
+To eliminate the need for expensive supervision, SUDS [50] and EmerNeRF [61] leverage the optical flow as decomposition guidance.
+PVG [5], VDG [22], S3Gaussian [16] and DeSiRe-GS [38] extend the self-supervised setting into 3DGS by a unified representation of both static and dynamic elements.
+However, existing self-supervised methods focus on the scene decomposition, overlooking the diverse motion patterns in street environments.
+This limitation becomes particularly pronounced when reconstructing objects moving at significantly different speeds.
+To tackle this challenge, we are the first to explore explicit motion modeling and introduce a plug-and-play Explicit Motion Decomposition (EMD) method.
+EMD enhances self-supervised approaches by capturing the continuous spectrum of motion, and it can be seamlessly extended to supervised methods.
 Figure 2: Overview of our Explicit Motion Decomposition (EMD) framework.
 Given input Gaussian primitives, our method processes them through two main components: (a) Motion-aware Feature Encoding, which combines spatial, temporal, and Gaussian-specific information to capture motion characteristics; and (b) Dual-scale Deformation Framework, which hierarchically models fast global motions and slow local deformations.
-The framework can be seamlessly integrated into both supervised (StreetGaussian) and self-supervised (S3Gaussian) approaches through our proposed integration strategies.
+The framework can be seamlessly integrated into existing supervised and self-supervised approaches.
 3 Preliminaries
 3.1 3D Gaussian Splatting
-3D Gaussian Splatting (3D-GS) (Kerbl et al.
-2023) proposes an explicit rendering approach to represent 3D scenes through a collection of 3D Gaussian primitives $\mathbb{G}=\left\{\left(\mu_{k},\Sigma_{k},\alpha_{k},\mathbf{c}_{k}\right)\right\}_{k=1}^{K}$, where $K$ is the total number of Gaussians.
+3D Gaussian Splatting (3DGS) [19] proposes an explicit rendering approach to represent 3D scenes through a collection of 3D Gaussian primitives $\mathbb{G}=\left\{\left(\mu_{k},\Sigma_{k},\alpha_{k},\mathbf{c}_{k}\right)\right\}_{k=1}^{K}$, where $K$ is the total number of Gaussians.
 Each Gaussian primitive represents a probabilistic distribution of density in 3D space, defined by its probability density function:
 $G_{k}(x)=e^{-\frac{1}{2}(x-\mu_{k})^{T}\Sigma_{k}^{-1}(x-\mu_{k})},$ (1)
-where $x\in\mathbb{R}^{3}$ represents any point in the 3D world space, $\mu_{k}\in\mathbb{R}^{3}$ and $\Sigma_{k}\in\mathbb{R}^{3\times 3}$ denote the mean position and covariance matrix in world space, respectively, where the covariance matrix determines the shape and orientation of the Gaussian, $\alpha_{k}\in[0,1]$ is the opacity, and $\mathbf{c}_{k}$ encodes the view-dependent color information.
-Rendering Process
-For rendering, each 3D Gaussian is projected onto the image plane, where the 3D mean $\mu_{k}$ is transformed to 2D mean $\mu_{k}^{2D}$, and the world space covariance matrix $\Sigma_{k}$ is transformed to screen space as $\Sigma_{k}^{\prime}=JW\Sigma_{k}W^{T}J^{T}$, with $W$ and $J$ being the viewing transformation and projective transformation Jacobian matrices, respectively.
-The final pixel color at screen space position $x$ is computed through front-to-back alpha compositing:
+where $x\in\mathbb{R}^{3}$ represents any point in the world space, $\mu_{k}\in\mathbb{R}^{3}$ and $\Sigma_{k}\in\mathbb{R}^{3\times 3}$ denote the mean position and covariance matrix in world space, respectively, where the covariance matrix determines the shape and orientation of the Gaussian, $\alpha_{k}\in[0,1]$ is the opacity, and $\mathbf{c}_{k}$ encodes the view-dependent color information with spherical harmonics.
+For rendering, each 3D Gaussian is projected onto the image plane, where the 3D mean $\mu_{k}$ is transformed to 2D mean $\mu_{k}^{2D}$, and the world space covariance matrix $\Sigma_{k}$ is transformed to screen space as $\Sigma_{k}^{\prime}=JW\Sigma_{k}W^{T}J^{T}$, with $W$ and $J$ being the viewing transformation and projective transformation Jacobian matrices.
+The pixel color at screen space position $x$ is computed through alpha blending:
 $C(x)=\sum_{k\in\mathcal{N}(\mathbf{x})}\mathbf{c}_{k}\alpha_{k}(x)\prod_{j=1}^{k-1}\left(1-\alpha_{j}(x)\right),$ (2)
-where $\alpha_{k}(x)=\alpha_{k}\exp\left(-\frac{1}{2}\left(x-\mu_{k}^{2D}\right)^{T}{\Sigma_{k}^{\prime}}^{-1}\left(x-\mu_{k}^{2D}\right)\right)$ represents the opacity contribution of the $k$-th Gaussian at pixel $x$, computed as the product of the Gaussian’s base opacity $\alpha_{k}$ and its projected 2D Gaussian evaluation.
+where $\alpha_{k}(x)=\alpha_{k}\exp\left(-\frac{1}{2}\left(x-\mu_{k}^{2D}\right)^{T}{\Sigma_{k}^{\prime}}^{-1}\left(x-\mu_{k}^{2D}\right)\right)$ represents the opacity contribution of the $k$-th Gaussian at pixel $x$.
 $\mathcal{N}(\mathbf{x})$ represents the set of indices of Gaussians intersecting pixel $x\in\mathbb{R}^{2}$.
@@
 $\Sigma_{k}=R(\mathbf{q}_{k})S(\mathbf{s}_{k})S(\mathbf{s}_{k})^{T}R(\mathbf{q}_{k})^{T},$ (3)
-where $R(\mathbf{q}_{k})$ is the rotation matrix defined by quaternion $\mathbf{q}_{k}$, and $S(\mathbf{s}_{k})$ is the diagonal scaling matrix defined by scaling vector $\mathbf{s}_{k}$.
-The view-dependent color $\mathbf{c}_{k}$ is encoded using spherical harmonics (SH) coefficients:
-$\mathbf{c}_{k}(\mathbf{d})=\sum_{l=0}^{L}\sum_{m=-l}^{l}k_{l,m}^{(k)}Y_{l,m}(\mathbf{d}),$ (4)
-where $\mathbf{d}$ is the viewing direction, $Y_{l,m}$ are the spherical harmonic basis functions, and $k_{l,m}^{(k)}$ are the corresponding coefficients for the $k$-th Gaussian.
+where $R(\mathbf{q}_{k})$ is the rotation matrix defined by $\mathbf{q}_{k}$, and $S(\mathbf{s}_{k})$ is the diagonal scaling matrix defined by $\mathbf{s}_{k}$.
 3.2 4D Gaussian Splatting
-4D Gaussian Splatting extends the static 3D-GS framework to handle dynamic scenes by incorporating temporal information.
+4D Gaussian Splatting (4DGS) [54, 63] extend the static 3DGS framework to handle dynamic scenes by incorporating temporal information.
 For a dynamic scene captured at different timestamps $t\in[0,T]$, each Gaussian primitive is now characterized by time-varying parameters: $\mathbb{G}(t)=\left\{\left(\mu_{k}(t),\Sigma_{k}(t),\alpha_{k}(t),\mathbf{c}_{k}(t)\right)\right\}_{k=1}^{K}$.
@@
 $G_{k}(x,t)=e^{-\frac{1}{2}(x-\mu_{k}(t))^{T}\Sigma_{k}(t)^{-1}(x-\mu_{k}(t))},$ (4)
-The temporal evolution of these parameters is typically modeled through one of two main approaches:
-Deformation-based Approach
-This approach models temporal changes by applying a deformation field to the base Gaussians. For each timestamp $t$, the position of each Gaussian is updated as:
+Then a deformation field is applied to Gaussians. For each timestamp $t$, the position of each Gaussian is updated as:
 $\mu_{k}(t)=\mu_{k}(0)+\Delta\mu_{k}(t),$ (5)
@@
 Similarly, other parameters such as rotation, scaling, and color can be modeled as temporal offsets from their initial states.
-Trajectory-based Approach
-Alternatively, the temporal evolution can be represented by explicitly modeling the continuous trajectory of each parameter.
-For instance, the position trajectory can be parameterized using a set of N control points $\{\mathbf{p}_{i}\}_{i=1}^{N}$ and basis functions $\{\phi_{i}(t)\}_{i=1}^{N}$:
-$\mu_{k}(t)=\sum_{i=1}^{N}\mathbf{p}_{i}\phi_{i}(t),$ (7)
-where the basis functions $\phi_{i}(t)$ can be B-splines or other temporal interpolation functions.
-The rendering process remains similar to static 3D-GS, but now Gaussian parameters are evaluated at the specific timestamp $t$ before projection and compositing.
 4 Methodology
-4D Gaussian Splatting has shown promising results in dynamic scene reconstruction.
+4DGS has shown promising results in dynamic scene reconstruction.
 However, modeling dynamic street scenes remains challenging due to diverse motion patterns.
-To better handle the varying motion patterns in street scenes, we propose Explicit Motion Decomposition (EMD), a plug-and-play module that can be seamlessly integrated into existing 4D Gaussian-based frameworks to enhance their capability in handling dynamic scenarios, as illustrated in Fig.
+To better handle the varying motion patterns in street scenes, we propose Explicit Motion Decomposition (EMD), a plug-and-play module that can be seamlessly integrated into existing street Gaussian Splatting methods to enhance their capability in handling dynamic scenarios, as illustrated in Fig.
 2.
@@
 To effectively handle the diverse motion patterns in street scenes, especially the distinct movements between vehicles and pedestrians, we propose a motion-aware deformation module that processes input Gaussian parameters through two key components: motion-aware feature encoding and dual-scale deformation prediction.
-Motion-aware Feature Encoding
+4.1.1 Motion-aware Feature Encoding
 Different types of objects in street scenes exhibit distinct motion characteristics.
@@
 $\mathbf{F}_{pos}(\mu)=[\mu,\{\sin(2^{i}\pi\mu),\cos(2^{i}\pi\mu)\}_{i=0}^{P-1}],$ (8)
-where $P$ is the number of frequency bands, enabling the network to capture both fine geometric details and global structures.
+where $P=10$ is the number of frequency bands, enabling the network to capture both fine geometric details and global structures.
 For temporal information, we design an adaptive temporal embedding function:
 $\mathbf{F}_{temp}(t)=\mathcal{T}(t,N(i))=\text{Interp}(\mathbf{W},t,N(i)),$ (9)
-where $\mathbf{W}\in\mathbb{R}^{N_{max}\times D}$ is a learnable embedding matrix that captures motion patterns, and $N(i)$ progressively increases from $N_{min}$ to $N_{max}$ temporal samples during training iteration $i$, allowing the model to gradually capture finer temporal dynamics.
-For Gaussian-specific features, we assign a learnable latent embedding $\mathbf{z}_{k}\in\mathbb{R}^{D}$ to each Gaussian $k$, where $\mathbf{F}_{gauss}=\mathbf{z}_{k}$, enabling the model to learn and represent individual motion characteristics.
-Table 1: Comparative performance of our framework and baseline approaches on the Waymo-NOTR dataset.
-The best performances are highlighted in bold, and the second-best are indicated with underlining.
-↑ indicates higher is better, while ↓ indicates lower is better.
-Scene Reconstruction Novel View Synthesis
-Dataset Methods Full Image Vehicle Full Image Vehicle
-PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑
-D32 EmerNeRF 28.16 0.806 0.228 24.32 0.682 25.14 0.747 0.313 23.49 0.660
-3DGS 28.47 0.876 0.136 23.26 0.716 25.14 0.813 0.165 20.48 0.753
-MARS 28.24 0.866 0.252 23.37 0.701 26.61 0.796 0.305 22.21 0.697
-S3Gaussian 30.69 0.900 0.121 26.23 0.804 26.62 0.824 0.159 22.61 0.681
-S3Gaussian+Ours 32.50 0.933 0.082 29.04 0.879 26.55 0.833 0.126 23.39 0.703
-Dual-scale Deformation Framework
+where $\mathbf{W}\in\mathbb{R}^{N_{max}\times D}$ is a learnable embedding matrix that captures motion patterns, and $N(i)$ progressively increases from $N_{min}=30$ to $N_{max}=150$ temporal samples during training iteration $i$, allowing the model to gradually capture finer temporal dynamics.
+$D=4$ is the temporal embedding dimension.
+During temporal embedding $\mathbf{W}$ optimization, we first downsample $\mathbf{W}$ through bilinear interpolation to create an intermediate embedding matrix of size $N(i)\times D$ at training iteration $i$, which can be formulated as:
+$N(i)=N_{min}+(N_{max}-N_{min})\cdot\min(i,T)/T,$ (10)
+where hyperparameter $T=25000$ controls the duration of progressive sampling refinement.
+Then, for each time stamp t, we perform grid sampling on this intermediate matrix with bilinear interpolation mode to obtain the corresponding temporal embedding vector $\mathbf{F}_{temp}(t)$.
+For Gaussian-specific features, we assign a learnable latent embedding $\mathbf{z}_{k}\in\mathbb{R}^{M}$ to each Gaussian $k$, where $\mathbf{F}_{gauss}=\mathbf{z}_{k}$ and $M=32$, enabling the model to represent individual motion characteristics.
+4.1.2 Dual-scale Deformation Framework
 Given the diverse motion patterns in street scenes, ranging from large vehicular movements to subtle pedestrian motions, we design a hierarchical deformation framework that can effectively handle both scales of motion:
 $\begin{split}\mathcal{D}(\mu,t)=&\mathcal{D}_{coarse}(\mathbf{F}_{aggr}(\mu,t))\\ &+\mathcal{D}_{fine}(\mathbf{F}_{aggr}(\mu+\Delta\mu_{coarse},t)),\end{split}$ (11)
-The final deformed parameters combine both coarse and fine scale predictions:
+The final deformed parameters combine both coarse and fine-scale predictions:
 $\mu_{t}$ $=\mu+\Delta\mu_{coarse}+\Delta\mu_{fine}$ (12)
@@
 Similar to position updates, other Gaussian parameters including opacity $\alpha_{t}$ and spherical harmonics coefficients $\mathbf{c}_{t}$ are also updated through this dual-scale framework.
-Figure 3: Qualitative comparison on the self-supervised setting between S3Gaussian and S3Gaussian+ours EMD.
-We also visualize the error maps between the rendered images and ground truth to provide further insights.
 4.2 Integration with Existing Frameworks
-As mentioned in the introduction, current approaches for street scene reconstruction generally fall into supervised and self-supervised paradigms.
-Having formalized our EMD framework, we now demonstrate its integration into representative methods from both paradigms: StreetGaussian for supervised learning and S3Gaussian for self-supervised learning.
-Self-supervised Integration: S3Gaussian
-For self-supervised scenarios, we enhance S3Gaussian’s architecture with our motion-aware features.
-While S3Gaussian originally employs a Multi-head Gaussian Decoder for deformation prediction, we augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and restructure its decoder into our dual-scale framework.
+Current approaches for street scene reconstruction generally fall into supervised and self-supervised paradigms.
+Having formalized our EMD framework, we now illustrate its integration into representative supervised methods.
+Then we further extend EMD into supervised approaches.
+4.2.1 Self-supervised Integration
+For self-supervised street Gaussian splatting methods, we enhance S3Gaussian [16] and DeSiRe-GS [38] with our motion-aware techniques.
+S3Gaussian originally employs a HexPlane to model the dynamic and static element decomposition and utilizes a Multi-head Gaussian Decoder for deformation prediction.
+We augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and restructure its decoder into our dual-scale framework.
 Specifically, both coarse and fine stages predict deformations in position ($\Delta\mu$), opacity ($\Delta\alpha$), and spherical harmonics coefficients ($\Delta\mathbf{c}$), enabling more precise motion modeling through hierarchical refinement.
-Supervised Integration: StreetGaussian
-For supervised settings, StreetGaussian provides a framework that represents dynamic objects through tracked vehicle poses and object-specific Gaussians.
+As for DeSiRe-GS built upon PVG [5], we also assign the learnable Gaussian embedding $\mathbf{z}_{k}$ into the framework.
+Then we apply the deformation for the position ($\Delta\mu$) and spherical harmonics coefficients ($\Delta\mathbf{c}$) with the proposed dual-scale deformation framework.
+We retain the same self-supervised deformation settings as those used in the baseline methods.
+Table 1: S3Gaussian comparison: results on the Waymo Open dataset for scene reconstruction and novel view synthesis.
+Scene Reconstruction Novel View Synthesis
+Dataset Methods Full Image Vehicle Full Image Vehicle
+PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑
+Waymo-D32 EmerNeRF [61] 28.16 0.806 0.228 24.32 0.682 25.14 0.747 0.313 23.49 0.660
+3DGS [19] 28.47 0.876 0.136 23.26 0.716 25.14 0.813 0.165 20.48 0.753
+MARS [55] 28.24 0.866 0.252 23.37 0.701 26.61 0.796 0.305 22.21 0.697
+S3Gaussian [16] 30.69 0.900 0.121 26.23 0.804 26.62 0.824 0.159 22.61 0.681
+S3Gaussian+Ours 32.50 0.933 0.082 29.04 0.879 26.55 0.833 0.126 23.39 0.703
+4.2.2 Supervised Extension
+For supervised settings, though existing methods refine tracked 3D bounding boxes to eliminate rendering errors, we extend EMD into the refinement process, illustrating the need for explicit motion modeling.
+We select StreetGaussian [60] and OmniRe [6] as our baselines.
+StreetGaussian provides a framework that represents dynamic objects through tracked vehicle poses and object-specific Gaussians.
 Each object is characterized by tracked poses $\{R_{t},T_{t}\}_{t=1}^{N_{t}}$ that transform object Gaussians from local coordinates ($\mu_{o},R_{o}$) to world coordinates ($\mu_{w},R_{w}$).
 To enhance its motion modeling capability, we also augment each Gaussian with our learnable embedding $\mathbf{z}_{k}$ and apply temporal embedding only within each object’s tracked duration, effectively capturing object-specific temporal dynamics.
-To address the challenge of noisy tracked poses while maintaining their geometric meaning, we incorporate our dual-scale framework into their pose optimization:
-$R^{\prime}_{t}$ $=R_{t}(\Delta R_{t}^{c}+\Delta R_{t}^{f})$ (14)
+Finally, we incorporate our dual-scale framework into the pose optimization:
+$R^{\prime}_{t}$ $=\Delta R_{t}^{f}\cdot\Delta R_{t}^{c}\cdot R_{t}$ (13)
 $T^{\prime}_{t}$ $=T_{t}+(\Delta T_{t}^{c}+\Delta T_{t}^{f}),$
 where $\Delta R_{t}^{c},\Delta T_{t}^{c}$ handle large pose corrections and $\Delta R_{t}^{f},\Delta T_{t}^{f}$ capture subtle adjustments.
-This decomposition allows the model to effectively correct tracking errors while preserving the physical meaning of the tracked poses.
-Similarly, we apply the dual-scale framework to appearance modeling, where spherical harmonics coefficients are refined through both coarse and fine stages, enabling more accurate dynamic appearance representation.
-We adopt the same loss function as S3Gaussian and StreetGaussian to train the entire pipeline. For further training details, please refer to the supplementary materials.
+Similarly, we apply the dual-scale framework to appearance modeling, where spherical harmonics coefficients are refined through both coarse and fine stages.
+For OmniRe, it further proposes non-rigid SMPL [27] nodes for human modeling.
+For the rigid nodes, we apply the same tracked box refinement in StreetGaussian into OmniRe.
+We further implement the dual-scale refinement to the SMPL model including pose parameters $\theta_{t}\in\mathbb{R}^{24\times 3\times 3}$ and shape parameters $\beta_{t}\in\mathbb{R}^{10}$:
+$\theta^{\prime}_{t}$ $=\Delta\theta_{t}^{f}\cdot\Delta\theta_{t}^{c}\cdot\theta_{t}$ (14)
+$\beta^{\prime}_{t}$ $=\beta_{t}+(\Delta\beta_{t}^{c}+\Delta\beta_{t}^{f}),$
+For training, in addition to using the same loss function as the baselines, we implement a local smoothness regularization for the learnable Gaussian embeddings.
+Inspired by  [29], this regularization encourages neighboring Gaussians $i$ and $j$ to have similar representations:
+$\mathcal{L}_{\mathbf{z}_{k}}=\frac{1}{d|\mathcal{U}|}\sum_{i\in\mathcal{U}}\sum_{j\in\text{KNN}_{i;d}}(e^{-\lambda_{w}\|\mu_{j}-\mu_{i}\|_{2}}\|\mathbf{z}_{k_{i}}-\mathbf{z}_{k_{j}}\|_{2}),$ (15)
+where hyperparameters $\lambda_{w}=2000$ and $d=20$.
+KNN means the k-nearest-neighbors algorithm.
+We also regularize the predicted coarse and fine deformations, constraining their values to remain close to zero.
+For further training details, please refer to the supplementary materials.
+Table 2: DeSiRe-GS comparison: results on the Waymo Open dataset and KITTI dataset for scene reconstruction and novel view synthesis.
+Method Waymo Open KITTI
+Scene Reconstruction Novel View Synthesis FPS Scene Reconstruction Novel View Synthesis FPS
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+S-NeRF [57] 19.67 0.528 0.387 19.22 0.515 0.400 0.0014 19.23 0.664 0.193 18.71 0.606 0.352 0.0075
+StreetSurf [13] 26.70 0.846 0.3717 23.78 0.822 0.401 0.097 24.14 0.819 0.257 22.48 0.763 0.304 0.37
+3DGS [19] 27.99 0.866 0.293 25.08 0.822 0.319 63 21.02 0.811 0.202 19.54 0.776 0.224 125
+NSG [34] 24.08 0.656 0.441 21.01 0.571 0.487 0.032 19.19 0.683 0.189 17.78 0.645 0.312 0.19
+Mars [55] 21.81 0.681 0.430 20.69 0.636 0.453 0.030 27.96 0.900 0.185 24.31 0.845 0.160 0.31
+SUDS [50] 28.83 0.805 0.317 25.36 0.783 0.384 0.008 28.83 0.917 0.147 26.07 0.797 0.131 0.29
+EmerNeRF [61] 28.11 0.786 0.373 25.92 0.763 0.384 0.053 26.95 0.828 0.218 25.24 0.801 0.237 0.28
+PVG [5] 32.46 0.910 0.229 28.11 0.849 0.279 50 32.83 0.937 0.070 27.43 0.896 0.114 59
+DeSiRe-GS [38] 33.61 0.919 0.204 29.75 0.878 0.213 36 33.94 0.949 0.040 28.87 0.901 0.106 41
+DeSiRe-GS + ours 34.15 0.948 0.130 29.91 0.880 0.190 32 34.13 0.954 0.040 29.05 0.904 0.094 32
+Figure 3: Visualization comparison on the self-supervised setting between S3Gaussian and S3Gaussian+ours EMD.
+We also visualize the error maps between the rendered images and ground truth to provide further insights.
+Please refer to appendix for more visualization.
+Figure 4: Motion deformation comparison between S3Gaussian + Ours and S3Gaussian. Please zoom in for more details.
 5 Experiments
-Figure 4: Motion deformation comparison between S3Gaussian + Ours and S3Gaussian. Please zoom in for more details.
-In this section, we first describe the datasets and evaluation metrics used in our experiments (Sec.
-5.1).
-Then, we demonstrate the experimental results of our approach under both self-supervised and supervised settings (Sec.
-5.2).
-Finally, we present comprehensive ablation studies to validate the effectiveness of individual components in our framework (Sec.
-5.3).
 5.1 Datasets and Metrics
-Dataset.
-We conduct extensive experiments on the Waymo Open Dataset (Sun et al.
-2020) to comprehensively evaluate our method.
-To validate the effectiveness of our plug-and-play module, we select representative state-of-the-art methods from both supervised and self-supervised paradigms: StreetGaussian (Yan et al.
-2024) and S3Gaussian (Huang et al.
-2024).
-For self-supervised evaluation, we use the dynamic32 (D32) split introduced by EmerNeRF (Yang et al.
-2023a), containing 32 sequences with vehicle motion, where each sequence consists of approximately 50-100 frames captured by three cameras under various conditions.
-For supervised evaluation, we follow the data split protocol established in StreetGaussian (Yan et al.
-2024) to enable direct comparison.
+Dataset and Baselines.
+We conduct extensive experiments on the Waymo Open Dataset [44] and KITTI Dataset [12] to comprehensively evaluate our method.
+For self-supervised setting, we select representative state-of-the-art methods including S3Gaussian [16] and DeSiRe-GS [38].
+To compare with S3Gaussian, we use the dynamic32 (D32) split with 3 frontal cameras of Waymo dataset introduced by EmerNeRF [61] and benchmark S3Gaussian + EMD with vanilla 3DGS [19] and MARS [55] on scene reconstruction and novel view synthesis.
+To compare with DeSiRe-GS which follows PVG [5], we use the same subset in PVG, including 4 Waymo scenes and 3 KITTI scenes.
+We implement baselines including S-NeRF [57], StreetSurf [13], NSG [35] and SUDS [50] on scene reconstruction and novel view synthesis.
+For the supervised setting, we select state-of-the-art approaches including StreetGaussian [60] and OmniRe [6].
+We select the same Waymo subset from StreetGaussian and OmniRe and conduct experiments on novel view synthesis.
+It should be noted that StreetGaussian only conducts experiments with one frontal camera in its published version.
+Thus, we further implement experiments on three frontal cameras for a side-by-side comparison.
 Evaluation Metrics.
-We employ comprehensive metrics to evaluate both reconstruction quality and novel view synthesis capability.
-For scene reconstruction, we use PSNR and SSIM to evaluate both full-scene and vehicle-specific reconstruction quality.
+Following previous protocols, we use PSNR and SSIM to evaluate the pixel-level reconstruction quality.
 Additionally, we compute LPIPS for perceptual quality assessment.
-Following previous protocols, we evaluate novel view synthesis on every 10th frame for self-supervised settings and every 4th frame for supervised settings.
-Figure 5: Qualitative ablation study results across three camera views from the Waymo dataset.
-(a) Our complete model achieves sharp and consistent reconstruction.
-(b) Removing Gaussian embedding $\mathcal{F}_{gauss}$ leads to blurry object boundaries.
-(c) Without temporal embedding $\mathcal{F}_{temp}$, results show motion artifacts.
-(d) Without coarse deformation $\mathcal{D}_{coarse}$, geometric consistency is lost.
-(e) Absence of fine deformation $\mathcal{D}_{fine}$ causes detail degradation.
+We also report FPS to access inference speed in the comparison between DeSiRe-GS, which refers to frames per second.
+In addition, we employ the FID metric [15, 65] for novel trajectory synthesis evaluation, which quantifies differences in feature distribution between rendered novel trajectory images and original trajectory images.
 5.2 Main Results
-Self-supervised Performance
+5.2.1 Self-supervised Performance
 Tab.
-1 presents the comparative results on the D32 split, where no 3D bounding box annotations were used.
+1 presents the comparative results on the Waymo-D32 split, where no 3D bounding box annotations were used.
 Our method significantly outperforms previous self-supervised approaches, achieving notable improvements in both full-scene and object-specific metrics.
-Specifically, we observe a substantial increase in PSNR for the full scene (32.50 vs.
-30.69) and for vehicle-specific metrics (PSNR: 29.04 vs.
-26.23) when compared to S3Gaussian.
-These results demonstrate the enhanced ability of our method to accurately model complex street scenes without explicit 3D box annotations.
-For novel view synthesis, our method continues to perform competitively, achieving a PSNR of 26.55, which reflects its robust generalization to previously unseen viewpoints.
+These results demonstrate the enhanced ability of our method to accurately model complex motions in street scenes without explicit 3D box annotations.
+S3Gaussian models the entire scene holistically but lacks the ability to capture the diverse motion patterns in dynamic street scenes, whereas EMD effectively compensates for this shortcoming.
+For novel view synthesis, our method continues to perform competitively, which reflects its robust generalization to previously unseen viewpoints.
+Tab.
+2 compares DeSiRe-GS with our proposed EMD.
+The results show that EMD achieves state-of-the-art performance across various experimental settings.
+DeSiRe-GS relies on dynamic masks to identify foreground objects, thereby overlooking motion patterns in street scenes.
+EMD boosts the rendering quality of DeSiRe-GS with only a slight reduction in inference speed.
+These quantitative results effectively demonstrate that EMD is well-suited for current self-supervised street Gaussian methods.
 In addition, we present a side-by-side visualization comparison between S3Gaussian and S3Gaussian+EMD in Fig.
@@
 This emphasizes the effectiveness of Explicit Motion Decomposition (EMD) in modeling the motion of dynamic objects, improving the overall decomposition and photorealistic rendering of street scenes.
+For more visualization videos, please watch the webpage Sec.B in the supplementary materials.
 We also present a visual comparison of motion between S3Gaussian and S3Gaussian+EMD.
@@
 In contrast, S3Gaussian treats the entire moving car as a single dynamic object, failing to produce clear synthesis results for the dynamic vehicles.
-Supervised Performance
-Table 2: Comparative performance of our framework and baseline approaches on the StreetGaussian dataset.
-The best performances are highlighted in bold.
-↑ indicates higher is better, while ↓ indicates lower is better.
+Table 3: StreetGaussin comparison: novel view synthesis results on the Waymo Open dataset with one and three frontal cameras.
 Methods Full Image Vehicle
 PSNR↑ SSIM↑ LPIPS↓ PSNR↑
-3DGS 29.64 0.918 0.117 21.25
-NSG 28.31 0.862 0.346 24.32
-MARS 29.75 0.886 0.264 26.54
-EmerNeRF 30.87 0.905 0.133 21.67
-StreetGaussian 34.61 0.938 0.079 30.23
-StreetGaussian + Ours 36.03 0.949 0.067 29.81
-To demonstrate the versatility of our framework, we also evaluate it in the supervised setting using the same scenes as StreetGaussian (Yan et al.
-2024) (Tab.
-2).
-When incorporating 3D bounding box supervision through our adaptive training scheme, our method achieves superior performance in full scene reconstruction, improving PSNR by 1.42dB (36.03 vs.
-34.61), SSIM by 1.1% (0.949 vs.
-0.938), and LPIPS by 15.2% (0.067 vs.
-0.079).
-These results demonstrate that our approach effectively enhances the overall scene reconstruction quality while maintaining competitive performance in dynamic object modeling.
-Please refer to the supplementary for visualization on the supervised setting.
-5.3 Ablation Studies
-Table 3: Ablation study on the D32 dataset showing the impact of different components in our framework.
-All variants are evaluated using the self-supervised setting.
-The best performances are highlighted in bold.
-↑ indicates higher is better, while ↓ indicates lower is better.
+One camera setting
+3DGS [19] 29.64 0.918 0.117 21.25
+NSG [34] 28.31 0.862 0.346 24.32
+MARS [55] 29.75 0.886 0.264 26.54
+EmerNeRF [61] 30.87 0.905 0.133 21.67
+StreetGaussian [60] 34.61 0.938 0.079 30.23
+StreetGaussian + Ours 35.41 0.942 0.070 30.96
+Three camera setting
+StreetGaussian 29.70 0.858 0.149 26.72
+StreetGaussian + Ours 29.84 0.869 0.145 26.83
+Table 4: OmniRe comparison: novel view synthesis results on the Waymo Open dataset with three frontal cameras.
+Methods Full Image Human Vehicle
+PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑
+EmerNeRF [61] 29.67 0.883 20.32 0.454 22.07 0.609
+3DGS [19] 25.57 0.906 16.62 0.387 16.00 0.407
+DeformGS [63] 27.72 0.922 17.30 0.426 18.91 0.530
+PVG [5] 30.19 0.919 21.30 0.567 22.28 0.679
+HUGS [67] 27.65 0.914 15.99 0.378 23.27 0.748
+StreetGS [60] 28.54 0.928 16.55 0.393 26.71 0.846
+OmniRe [6] 32.57 0.942 24.36 0.727 27.57 0.858
+OmniRe + Ours 33.89 0.958 25.97 0.742 27.82 0.859
+5.2.2 Supervised Performance
+To demonstrate the flexibility of EMD, we further extend it to enhance motion modeling in supervised methods.
+Table 3 and Table 4 show novel view synthesis comparisons between StreetGaussian and OmniRe, respectively.
+While these supervised methods employ 3D box refinement to mitigate tracking errors, the results indicate that EMD serves as a valuable complement to this process.
+In the OmniRe comparison, EMD also benefits non-rigid objects such as humans, whose motion patterns differ significantly from those of vehicles.
+Due to the space limitation, please refer to the supplementary for visualization on the supervised setting.
+Table 5: Novel trajectory synthesis on Waymo dataset.
+Method FID↓
+0.5m 1.0m 1.5m
+S3Gaussian 83.48 110.11 134.38
+S3Gaussian + ours 45.11 70.26 90.20
+Figure 5: Visualization for novel trajectory synthesis (0.5m offset). Please watch the webpage in the supplementary materials for more results.
+5.3 Novel Trajectory Synthesis
+Previous novel view synthesis evaluations are limited to interpolated views along the original camera trajectory, which fail to assess the exact simulation performance of the model.
+Therefore, we compare the FID performance of self-supervised methods and our approach under novel trajectory synthesis.
+Specifically, we shift the original camera trajectory to the left and right by different offsets (0.5 m, 1.0 m, 1.5 m) and render images along these new trajectories.
+The results are presented in Tab.
+5 and one visualization sample is shown in Fig.
+5.
+These findings demonstrate that EMD enhances the accurate modeling of dynamic objects by learning diverse motion patterns, which benefits lane change scenarios.
+For novel trajectory synthesis videos, please refer to the supplementary materials.
+Table 6: Ablation study on the 32 dynamic scenes Waymo subset showing the impact of different components in our framework. All variants are evaluated in self-supervised scene reconstruction.
 Variant Full Image Vehicle
@@
 w/o Fine Deformation 32.45 0.931 0.118 28.80
+Figure 6: Qualitative ablation study results across three camera views from the Waymo dataset.
+(a) Our complete model achieves sharp and consistent reconstruction.
+(b) Removing Gaussian embedding $\mathcal{F}_{gauss}$ leads to blurry object boundaries.
+(c) Without temporal embedding $\mathcal{F}_{temp}$, results show motion artifacts.
+(d) Without coarse deformation $\mathcal{D}_{coarse}$, geometric consistency is lost.
+(e) Absence of fine deformation $\mathcal{D}_{fine}$ causes detail degradation.
+5.4 Ablation Studies
 To assess the contribution of each component in our framework, we perform comprehensive ablation studies, as shown in Tab.
 6 and Fig.
-5.
+6.
 The results reveal the critical role of the Gaussian embedding, as its removal leads to the performance drop, with a reduction of 0.29 PSNR (32.21 vs.
@@
 32.50), underscoring its importance in modeling the temporal variation of object motion over time.
-Both the coarse and fine deformation components are integral to the final performance, with their removal leading to considerable performance degradation.
+Both the coarse and fine deformation components are integral to the final performance, with their removal leading to performance degradation.
 Specifically, excluding the coarse deformation component causes a significant 3.10 PSNR drop (29.40 vs.
@@
 The ablation study demonstrates that our proposed motion-aware feature encoding and dual-scale deformation effectively model dynamic objects with varying motion speeds, which is crucial for enhancing the reconstruction quality of existing street Gaussians.
-6 Limitation
+6 Conclusion
+In this paper, we present EMD, the first plug-and-play module that effectively handles varying motion speeds in street scene reconstruction.
+By introducing motion-aware feature encoding and dual-scale deformation modeling, our approach successfully captures complex motion patterns in real-world scenarios.
+Comprehensive experiments on the Waymo and KITTI datasets demonstrate that EMD achieves the state-of-the-art novel view synthesis quality for self-supervised frameworks.
+EMD can be extended into supervised street Gaussian splatting methods, which opens up new possibilities for high-quality driving scene reconstruction.
+Acknowledgments.
+This work was supported by the National Science and Technology Major Project (No. 2022ZD0117800).
+Supplementary Material
+Appendix A Overview
+The supplementary material includes the subsequent components.
+  • Additional Visualization Videos
+  • Implementation Details
+    • Training Schemes
+    • Training Details
+    • Parameters and Efficiency
+  • Parameter Sensitivity
+Appendix B Additional Visualization Videos
+Please double-click the “Demo Webpage-Please wait until loaded.html" file and open it in your browser. This offline webpage contains videos covering the following experiments:
+  • Self-supervided Comparison
+  • Box Supervised Comparison
+  • Novel Trajectory Synthesis
+  • Temporal Embedding Evolution
+Due to the numerous videos, please wait for the webpage until loaded.
+Appendix C Implementation Details
+C.1 Training Schemes
+LiDAR Prior Initialization.
+To initialize the positions of the 3D Gaussians, we leverage the LiDAR point cloud captured by the vehicle instead of using the original SFM [42] point cloud to provide a better geometric structure.
+To reduce model size, we also downsample the entire point cloud by voxelizing it and filtering out points outside the image.
+For colors, we initialize them randomly.
+Optimization Objective.
+Following Street Gaussian, we introduce the sky supervision loss $L_{sky}$ into the original loss function proposed by S3Gaussian.
+Subsequently, we get a composed training loss function which can impose various constraints to our model.
+$\begin{split}\mathcal{L}=\mathcal{L}_{color}&+\lambda_{depth}\mathcal{L}_{depth}+\lambda_{feat}\mathcal{L}_{feat}\\ &+\lambda_{tv}\mathcal{L}_{tv}+\lambda_{sky}\mathcal{L}_{sky}+\lambda_{reg}\mathcal{L}_{reg}\end{split}$ (16)
+Here, $\mathcal{L}_{depth}$ is the mean square error (MSE) loss between the rendered depth map and the estimated depth map from the LiDAR point cloud, which aids in supervising the expected position of 3D Gaussians.
+$\mathcal{L}_{feat}$ is also the L2 loss of semantic features to reduce the gap between both planes.
+$\mathcal{L}_{tv}$ is a total-variational loss based on grids to make rendered objects smoother.
+$\mathcal{L}_{color}$ is the main loss to give constraints to the reconstruction process formulated by:
+$\mathcal{L}_{color}=\mathcal{L}_{rgb}+\lambda_{ssim}\mathcal{L}_{sim}$ (17)
+Furthermore, $\mathcal{L}_{reg}$ is organized as:
+$\mathcal{L}_{reg}=\mathcal{L}_{z_{k}}+\mathcal{L}_{\Delta}$ (18)
+where $\mathcal{L}_{z_{k}}$ is local smoothness regularization for Gaussian embeddings in the method section.
+$\mathcal{L}_{\Delta}$ represents a combination of regularization for coarse and fine deformations, restricting their values near zero.
+We also detail the coefficients for loss in Tab.
+7.
+Table 7: Loss function coefficients
+$\lambda_{depth}$ $\lambda_{feat}$ $\lambda_{feat}$ $\lambda_{tv}$ $\lambda_{sky}$ $\lambda_{reg}$
+0.5 0.1 0.1 0.1 0.1 0.01
+Table 8: Parameter sensitivity analysis on the D32 dataset, highlighting the effect of varying the dimensions of Gaussian embeddings $\mathbf{z}_{k}$ and temporal embeddings $\mathbf{z}_{w}$.
+All experiments are conducted in the self-supervised setting.
+Best performances are highlighted in bold.
+↑ indicates higher is better, while ↓ indicates lower is better.
+We also include the changes in model parameters relative to the adopted setting.
+$\mathbf{z}_{k}$/$\mathbf{z}_{w}$ Parameters Full Image Vehicle
+PSNR↑ SSIM↑ LPIPS↓ PSNR↑
+32/4 / 32.50 0.933 0.082 29.04
+128/4 +14400 32.22 0.925 0.086 29.05
+8/4 -3600 31.25 0.910 0.128 27.75
+32/4 / 32.50 0.933 0.082 29.04
+32/16 +14.42M 32.38 0.930 0.081 29.01
+32/1 -3.60M 30.55 0.898 0.136 27.04
+C.2 Training Details
+For S3Gaussian, we train the entire pipeline for 50,000 iterations using the Adam optimizer.
+Following the original S3Gaussian setup, we perform a warm-up phase for each scene, employing 5,000 iterations to train a coarse representation using vanilla 3D Gaussians.
+After this warm-up phase, we integrate the proposed dual-scale deformation network, which is jointly optimized with the HexPlane.
+To implement a coarse-to-fine training strategy, temporal embeddings $N(i)$ are progressively increased from $N_{min}$ to $N_{max}$ in 20,000 iterations, allowing for the gradual motion modeling of objects.
+Since S3Gaussian is evaluated on 50 frames per clip for each scene, we ensure a fair comparison by conducting all self-supervised validation experiments on the first clip of 32 dynamic scenes.
+Other configurations, including the detailed setup of the HexPlane and learning rates, are kept consistent with the S3Gaussian.
+For StreetGaussian, the entire method is trained for 30,000 iterations on a subset of eight selected scenes from the StreetGaussian dataset.
+Unlike the self-supervised method, we bind the proposed EMD to the vehicle Gaussians in each scene.
+Temporal embeddings are applied based on the time each vehicle appears within the scene.
+All other settings, including detailed configurations, remain consistent with those described in StreetGaussian.
+All experiments are conducted on a single NVIDIA A800 GPU.
+Appendix D Parameter Sensitivity
+To analyze the sensitivity of model performance to the dimensions of Gaussian embeddings $\mathbf{z}_{k}$ and temporal embeddings $\mathbf{z}_{w}$ (derived from the learnable embedding matrix $\mathbf{W}$), we conduct experiments by varying these dimensions.
+In the original setup, $\mathbf{z}_{k}$ is set to 32 and $\mathbf{z}_{w}$ to 4.
+Tab.
+8 summarizes the results, demonstrating how these changes influence performance under the self-supervised setting.
+The results reveal that reducing the embedding dimensions leads to significant performance degradation.
+This is primarily due to the reduced capacity to effectively model the motion of dynamic objects, which is crucial for high-quality reconstruction.
+On the other hand, increasing the embedding dimensions offers only marginal performance improvements.
+However, due to the large number of Gaussians in the driving scenes, the higher embedding dimensions result in a substantial increase in the total number of model parameters, leading to a higher computational cost.
+These findings highlight the trade-off between embedding dimension size and overall model efficiency.
+While lower dimensions compromise the ability to capture dynamic motion, higher dimensions introduce considerable overhead without proportional gains in performance.
+The adopted embedding configuration achieves a good balance, maintaining strong performance while keeping the parameter count manageable.
+Appendix E Limitation
 Although EMD effectively addresses the challenge of modeling dynamic objects with varying speeds by incorporating learnable embeddings, some limitations remain.
@@
 In future work, we plan to explore the possibility of developing a plug-and-play technique to enhance lighting effects in existing methods.
-7 Conclusion
-In this paper, we present EMD, a plug-and-play module that effectively handles varying motion speeds in street scene reconstruction.
-By introducing motion-aware feature encoding and dual-scale deformation modeling, our approach successfully captures the continuous spectrum of motion patterns inherent in real-world scenarios.
-Comprehensive experiments on the Waymo Open Dataset demonstrate that EMD significantly improves reconstruction quality when integrated with both supervised and self-supervised frameworks.
-Our method substantially improves full scene reconstruction (+1.81 PSNR) and vehicle-specific regions (+2.81 PSNR), validating the effectiveness of explicit motion modeling.
-We believe our work opens up new possibilities for high-quality dynamic scene reconstruction in autonomous driving applications, and the plug-and-play nature of our approach makes it readily applicable to future developments in neural scene representations.